What value resistor will discharge a 2.40 μF capacitor to 20.0% of its initial charge in 3.00 ms ?
Answer #1

tIRc we Knoton H Vo e tRc. elt)0e Guen alt 0.2 O0 C 40ft =3.00mg 0.a etIRc un (0-a) Rc t Rc - +1.609 Rt -609 xC 3x103 -609 x
Answer #2

SOLUTION :


We use the following formula :


Q = Q0 e^(- t / R C) 

=> Q / Q0 = e^ ( - t / RC)


where, 


Q / Q0 = 0.20 (capacitor is discharged to 20% of initial value)

t = time taken in secs for discharge to 0.20 Q0  is 3 ms = 3^10^(- 3) sec.

R = Resistance  in Ωs  to be determined .

C = Capacitor capacitance = 2.40 µF= 2.40*10^(-6) F .


Hence,


0.20 =  e^(- 3*10^(-3) / (R * 2.40*10^(-6)) 

=> 0.20 = e^(- 1250 / R)

Taking natural logarithm :

=> ln(0.2) = - 1250 / R

=> R = - 1250 / ln(0.2) 

=> R = 776.67 Ω   approx. 


=> Resistor of 776.67  Ω approx. (ANSWER).

answered by: Tulsiram Garg